OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b... (-7,7); 7x=9y + 5 y = ?

OpenStudy (aroub):

parallel=same slope y=mx+b , m= 7/9 , (-7,7) 7=7/9(-7)+b 7=-49/7+b -49/7+49/7=b 0=b y= 7/9x

OpenStudy (anonymous):

7/9x

OpenStudy (aroub):

7/9 is the slope and the y-intercept is 0

OpenStudy (anonymous):

Is 7/9 the equation of the line? Sorry to sound so freaking dumb.

OpenStudy (anonymous):

I mean 7/9x sorry

OpenStudy (aroub):

well the equation of the line is y=7/9x / lol its okay!! :)

OpenStudy (anonymous):

thanks!

OpenStudy (aroub):

anytime!! :D